how many fl. oz. of a 20% acid solution must be mixed with 9 fl. oz. of a 40% acid solution to make a 35% solution
Answer by TimothyLamb(4379)

x = oz of 20% acid
---
0.20x + 0.40*9 = 0.35*(x + 9)
0.20x + 3.6 = 0.35x + 3.15
0.35x - 0.20x = 3.6 - 3.15
0.15x = 0.45
x = 0.45/0.15
x = 3
---
answer:
3 fl.oz. of 20% acid
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---